# Airway Labeling Meets Clinical Applications: Reflecting Topology Consistency and Abnormality via Learnable Attentions

### v1

@@ Abstract
Purpose: Accurate airway anatomical labeling is crucial for clinicians to identify and navigate complex bronchial structures during bronchoscopy. Automatic airway anatomical labeling is challenging due to significant individual variability and anatomical variations. Previous methods are prone to generate inconsistent predictions, which is harmful for preoperative planning and intraoperative navigation. This paper aims to address these challenges by proposing a novel method that enhances topological consistency and improves the detection of abnormal airway branches.
Methods: We propose a novel approach incorporating two modules: the Soft Subtree Consistency (SSC) and the Abnormal Branch Saliency (ABS). The SSC module constructs a soft subtree to capture clinically relevant topological relationships, allowing for flexible feature aggregation within and across subtrees. The ABS module facilitates the interaction between node features and prototypes to distinguish abnormal branches, preventing the erroneous aggregation of features between normal and abnormal nodes.
Results: Evaluated on a challenging dataset characterized by severe airway distortion and atrophy, our method achieves superior performance compared to state-of-the-art approaches. Specifically, it attains a 91.4% accuracy at the segmental level and an 83.7% accuracy at the subsegmental level, representing a 1.4% increase in subsegmental accuracy and a 3.1% increase in topological consistency. Notably, the method demonstrates reliable performance in cases with disease-induced airway deformities, ensuring consistent and accurate labeling.
Conclusion: The enhanced topological consistency and robust identification of abnormal branches provided by our method offers a reliable solution for airway labeling, with significant potential to improve both the effectiveness and safety of bronchoscopy procedures in clinical practice.
@@
Airway anatomical labeling, Structural prior, Anomaly detection
@@ 1 Introduction
Bronchial anatomical labeling enables clinicians to precisely identify the bronchus or subsegment involved with the pathological lesion, thereby facilitating the optimal planning of the surgical approach while minimizing injury to surrounding healthy tissues [1, 2]. When integrated with advanced navigation technologies, bronchial anatomical labeling supports the precise guidance of surgical instruments by robotic or navigational systems, ensuring accurate execution of procedures within the intricate pulmonary anatomy [3]. Furthermore, bronchial anatomical labeling also assists clinicians in better assessing individual anatomical variations in each patient and developing personalized surgical plans, particularly for patients with significant anatomical anomalies.
@@
Manual annotation of bronchial anatomical labeling by medical experts is time-consuming and requires a high level of expertise. Different from the binary segmentation tasks [4], this task assigns the branches with specific class labels according to the anatomy. Therefore, automated labeling algorithms have emerged to alleviate the workload of clinicians. However, automatic bronchial anatomical labeling among different patients can be challenging from two aspects. Firstly, individual variability leads to substantial feature overlap, further aggravated by disease-induced airway deformities. Traditional algorithms rely on predefined rules [1, 5] or association graph search [6, 7], failing to accurately label distal branches. Recently, deep learning methods [8, 9, 10] have demonstrated superior performance on this task. To address individual differences, a hyper-graph model was introduced by [8] to encode subtree representations through hyperedges. However, it relies heavily on manually designed hypergraphs and is sensitive to hyperparameters. Xie et al. [10] encoded deep point-graph representation for analyzing 3D pulmonary tree structures, but only achieved segmental-level classification. AirwayFormer [9] was proposed to aggregate features using a self-attention mechanism and neighborhood encoding. Although these approaches achieved promising labeling results, limited attention has been paid to airway deformity such as atrophy, compression, and dilatation, resulting in failure in deformed airways as shown in Fig. 1(b). Secondly, the bronchial tree exhibits considerable morphological variability, with many anatomical variations beyond established nomenclatures. These abnormal branches without certain locations and orientations lack consistent anatomical descriptions, which complicates identification and impacts the classification of other branches. However, current methods often overlook these variant branches. Huang et al. [11] treated abnormal bronchus as a separate category, while the absence of anatomical standards made it challenging to categorize them equally with other classes as illustrated in Fig. 1(c).
Figure 1: (a) illustrates the preoperative planning path for bronchoscopy. (b) and (c) shows that two main challenges will reduce clinical usability.
@@
To alleviate the aforementioned challenges, we focus on the topological consistency for bronchial anatomical labeling. Specifically, we propose the Soft Subtree Consistency (SSC) module to tackle the first challenge, SSC constructs a soft subtree via feature interaction and topological refinement. Segmental labels serve as the criterion for soft subtrees, resulting in more flexible and discriminative feature aggregation within and across subtrees. To deal with the second challenge, we propose an Abnormal Branch Saliency (ABS) module, which compares node features and prototypes to predict anomaly scores, distinguishing abnormal nodes. And a soft mask derived from anomaly scores is utilized to prevent feature aggregation between normal and abnormal nodes for a smoother regularization.
@@
The proposed method is evaluated on the AIIB23 [12] dataset, which is characterized by severe airway distortion and atrophy. Extensive experiments demonstrate that our method achieves superior performance compared to other state-of-the-art methods, achieving 91.4% accuracy in segmental, 83.7% accuracy in sub-segmental with a 1.4% increase, 96.8% consistency with a 3.1% increase, achieving consistent results in disease-induced airway deformities.
@@ 2 Method
Figure 2: Overview of the proposed method: The airway labeling method comprises a Soft Subtree Consistency (SSC) module, which generates a soft subtree probability map, and an Abnormal Branch Saliency (ABS) module, which produces an abnormality score prediction. Both of them are used to inform attention interaction at the subsegmental level.
Each airway branch is represented as a node in the graph, with attributes designed to reflect its location, morphology, and structure. An overview of the proposed method is illustrated in Fig. 2. We propose two modules to improve the consistency of prediction: (1) Soft Subtree Consistency (SSC) module (detailed in Section 2.2) softly encodes the relationship between subtrees. (2) Abnormal Branch Saliency (ABS) module (detailed in Section 2.3) distinguishes between normal nodes and abnormal nodes.
@@ 2.1 Overall Framework
We adopt a U-shaped framework that progressively improves classification accuracy by incorporating hierarchical relationships. This framework comprises two stages: an independent stage ($1st$ stage) and a guided refinement stage ($2nd$ stage). The latter employs a fine-to-coarse guidance mechanism, using finer level features from the independent stage to refine coarser-level predictions, thereby enhancing overall labeling performance.
@@
$\tilde{G}_{m}^{i}=\mathcal{F}(q=k=v=X_{m}^{i}),$
$G_{m}^{i}=\left\{\begin{array}[]{lll}\mathcal{F}(q=\tilde{G}_{m}^{i},k=v=G_{m+1}^{i-1})&&\text{for }m\in\{lob,seg\}\text{ and }i=2\\ \tilde{G}_{m}^{i}&&\text{otherwise}\end{array}\right..$ (3)
Here, the hierarchical level $m\in\mathcal{H}$, where the ordered set =$\mathcal{H}=\{lob,seg,sub\}$ corresponding to lobar, segmental, and subsegmental levels respectively. $\mathcal{F}$ denotes the function of Transformer architecture. $X_{m}^{i}\in\mathbb{R}^{N\times d}$ and $G_{m}^{i}\in\mathbb{R}^{N\times d}$ represent the input and output features of the m-th transformer block in the i-th stage, where $N$ denotes the number of nodes and $d$ is the feature dimension. For simplicity, we take the guided refinement stage for an example, omitting the index $i$ in subsequent notation.
@@ 2.2 Soft Subtree Consistency
Due to significant inter-individual variability, relying solely on node representations for classification can lead to substantial feature overlap. To address this challenge, we propose the Soft Subtree Consistency (SSC) module, which encodes subtree representations. Unlike hard-subtree techniques that use the number of descendants as a criterion for subtree construction, we utilize segmental labels to define subtrees that convey clinically relevant anatomical meaning. The soft subtree map $M_{t}\in\mathbb{R}^{N\times N}$ is introduced, where $M_{t}(i,j)$ represents the probability that nodes $v_{i}$ and $v_{j}$ belong to the same subtree. The ground truth labels for $M_{t}$ are defined as,
$M_{t}(i,j)=\left\{\begin{array}[]{lll}1&&Y_{seg}(i)=Y_{seg}(j)\\ 0&&Y_{seg}(i)\neq Y_{seg}(j)\\ \end{array}\right.,$ (7)
where $Y_{seg}\in\mathbb{R}^{N}$ denotes the segmental level label.
As segmental labels define subtrees, we utilize segmental feature $G_{seg}=(g_{1},g_{2},\dots,g_{N})^{T}\in\mathbb{R}^{N\times d}$ to infer the soft subtree map (as shown in Fig. 2). Taking node $v_{i}$ as an example, its feature $g_{i}$ interacts with the features of all the nodes, then processed by several $1\times 1$ convolutional layers to reduce dimensionality. Repeating this process for all nodes yields the soft subtree map $\bar{M}_{t}$. However, a hard subtree comprises a node and all its descendants. Similarly, to reflect this in our soft subtree, the probability of nodes $v_{i}$ and $v_{j}$ belonging to the same subtree should be increased if $v_{j}$ is a descendant of $v_{i}$. Hence, the soft topological refinement is defined as below:
$\hat{M}_{t}=M_{D}\odot(\bar{M}_{t}+K(1-\bar{M}_{t}))+(1-M_{D})\odot\bar{M}_{t},$
$M_{D}(i,j)=\left\{\begin{array}[]{lll}1&&v_{j}\in\mathcal{D}(v_{i})\\ 0&&v_{j}\notin\mathcal{D}(v_{i})\\ \end{array},\right.$ (10)
where $K\in[0,1]^{N\times 1}$ is an learnable parameter. $\mathcal{D}(v_{i})$ denotes the set of all descendant nodes of node $v_{i}$.
@@
According to the hierarchical nomenclature, the assignment of subsegmental labels is determined by their relative relationships to branches within a segment. The soft subtree map constrains feature aggregation within and across subtrees at the subsegmental level as follows,
@@
An example of the mask is shown in Fig. 3. As illustrated, the positions and orientations of LB4 branches in this variant differ significantly from the typical pattern. LB4 exhibits greater positional and orientational similarity to LB3 than LB5, resulting in substantial feature overlap between LB3 and LB4. The mask generated by the SSC module limits feature aggregation between different segments, thereby alleviating feature overlap.
@@ 2.3 Abnormal Branch Saliency
The position, orientation, and shape of abnormal branches are uncertain, with significant individual differences. The primary criterion for identifying an abnormal branch is its inability to be classified into any category within the nomenclature. Thus, we propose an Abnormal Branch Saliency (ABS) that compares node features with class representations to detect anomalies. First, the class representation is obtained through the segmental features $G_{seg}\in\mathbb{R}^{N\times d}$ and prediction logits $Z_{seg}\in\mathbb{R}^{N\times C_{seg}}$, where $C_{seg}$ denotes the number of segment categories to obtain:
$P_{seg}$ $=softmax(Z_{seg}),$ (12)
$H$ $=((P_{seg})^{\alpha})^{T}G_{seg},$ (13)
$P_{seg}\in[0,1]^{N\times C_{seg}}$ denotes prediction probabilities. $H\in\mathbb{R}^{C_{seg}\times d}$ are representations for segmental categories. Then they are refined through a transformer for improvement:
@@
Subsequent operations follow a similar process as the SSC module. Each pair of feature and prototypes is concatenated along the depth dimension, forming the $\phi\in\mathbb{R}^{N\times C_{seg}\times 2d}$. $\phi$ is processed through several $1\times 1$ convolutional layers to reduce the depth dimension, followed by the Multilayer Perceptron (MLP).
Finally, the ABS module generates an anomaly score $\hat{Y}_{a}\in[0,1]^{N\times 1}$, indicating the probability of each node being an abnormal branch. Due to the distinctive characteristics of abnormal branches, it is too absolute and arbitrary to use this anomaly score to directly classify nodes as a category equally with other classes. Instead, we generate a soft mask $M_{a}\in\mathbb{R}^{N\times N}$ for smoother regularization, which prevents feature aggregation between normal and abnormal nodes:
@@
Mask $\hat{M}_{a}$ guides abnormal nodes away from normal nodes in the feature space, aiding in more accurate predictions for normal nodes. Additionally, the abnormality score highlights potential abnormal branches and assists doctors in understanding the airway structure. Combining the two soft masks, $\hat{M}_{t}\odot\hat{M}_{a}$ is applied to the attention interaction of subsegmental Transformer Block,
@@
For classification tasks, cross-entropy loss with label smoothing is used, excluding abnormal nodes from the loss calculation. Binary cross-entropy loss is employed for both soft subtree map $\hat{M}_{t}$ prediction and abnormal score prediction $\hat{Y}_{a}$. The total loss function is formulated as:
$\begin{split}\mathcal{L}=\sum_{i=1}^{2}\gamma_{i}\Bigl{(}&\sum_{m\in\mathcal{H}}\alpha_{m}\mathcal{L}_{CE}(\hat{Y}_{m}^{i},Y_{m})\\ &+\beta_{1}\mathcal{L}_{BCE}(\hat{M}_{t}^{i},M_{t})+\beta_{2}\mathcal{L}_{BCE}(\hat{Y}_{a}^{i},Y_{a})\Bigr{)}\end{split},$ (17)

### v2

# Reflecting Topology Consistency and Abnormality via Learnable Attentions for Airway Labeling

@@ Abstract
Purpose: Accurate airway anatomical labeling is crucial for clinicians to identify and navigate complex bronchial structures during bronchoscopy. Automatic airway labeling is challenging due to significant anatomical variations. Previous methods are prone to generate inconsistent predictions,hindering preoperative planning and intraoperative navigation. This paper aims to enhance topological consistency and improve the detection of abnormal airway branches.
Methods: We propose a transformer-based framework incorporating two modules: the Soft Subtree Consistency (SSC) and the Abnormal Branch Saliency (ABS). The SSC module constructs a soft subtree to capture clinically relevant topological relationships, allowing for flexible feature aggregation within and across subtrees. The ABS module facilitates interaction between node features and prototypes to distinguish abnormal branches, preventing the erroneous features aggregation between normal and abnormal nodes.
Results: Evaluated on a challenging dataset characterized by severe airway deformities, our method achieves superior performance compared to state-of-the-art approaches. Specifically, it attains an 83.7% subsegmental accuracy, along with a 3.1% increase in segmental subtree consistency, a 45.2% increase in abnormal branch recall. Notably, the method demonstrates robust performance in cases with airway deformities, ensuring consistent and accurate labeling.
Conclusion: The enhanced topological consistency and robust identification of abnormal branches provided by our method offer an accurate and robust solution for airway labeling, with potential to improve the precision and safety of bronchoscopy procedures.  [^1] The code is publicly available at Reflecting-Topology-Consistency-and-Abnormality-via-Learnable-Attentions
@@
Airway anatomical labeling, Structural prior, Anomaly detection, Transformer
@@ 1 Introduction
Bronchial anatomical labeling enables clinicians to precisely identify the bronchus or subsegment involved with the pathological lesion, thereby facilitating the optimal planning of the surgical approach while minimizing injury to surrounding healthy tissues [1]. When integrated with advanced navigation technologies, the hierarchical bronchial labeling (lobar, segmental, and subsegmental levels) supports the precise guidance of surgical instruments by robotic or navigational systems(as illustrated in Fig. 1(a) and (b)). This ensures accurate execution of procedures within the intricate pulmonary anatomy [2]. Accurate anatomical labeling ensures reliable navigation trajectories for bronchoscopy; while labeling errors can lead to failed interventions, as shown in Fig. 1(c) and (d).
In addition, bronchial anatomical labeling allows for a comprehensive assessment of resectability and improved identification of complicated anatomical variations for segmentectomy [3]. Furthermore, its applications extend beyond procedural applications to include diagnostic advancements [4]. These features serve as robust biomarkers for clinical diagnosis and prognosis.
These clinical applications focus on local segments and subsegments surrounding the lesions, underscoring the importance of ensuring the precision and completeness of each anatomical region. As shown in Fig. 1, these regions can naturally be represented as anatomical subtrees within the airway structure. Unlike branch-wise correctness, anatomical subtree-wise consistency reflects the integrity of the entire anatomical structure, as demonstrated in Fig. 1(e), which is critical to deliver precise and clinically meaningful outcomes.
@@
Manual annotation of bronchial anatomical labeling by medical experts is time-consuming and requires a high level of expertise. Different from the binary segmentation tasks [5], this task assigns the branches with specific class labels according to the anatomy. Therefore, automated labeling algorithms have emerged to alleviate the workload of clinicians. However, automatic bronchial anatomical labeling among different patients can be challenging from two aspects. Firstly, individual variability leads to substantial feature overlap, further aggravated by disease-induced airway deformities. Traditional algorithms rely on predefined rules [1, 6] or association graph search [7, 8], failing to accurately label distal branches. Recently, deep learning methods [9, 10, 11] have demonstrated superior performance on this task. To address individual differences, Yu et al. [9] introduced a hypergraph model to encode subtree representations through hyperedges. However, it relies heavily on manually designed hypergraphs and is sensitive to hyperparameters. Xie et al. [11] encoded deep point-graph representation for analyzing 3D pulmonary tree structures, but only achieved segmental level classification. AirwayFormer [10] was proposed to aggregate features using a self-attention mechanism and neighborhood encoding. Although these approaches achieved promising labeling results, limited attention has been paid to airway deformity such as atrophy, compression, and dilatation, resulting in failure in deformed airways as shown in Fig. 1(f). Secondly, the bronchial tree exhibits considerable morphological variability, with many anatomical variations beyond established nomenclatures. These abnormal branches without certain locations and orientations lack consistent anatomical descriptions, which complicates identification and impacts the classification of other branches. However, current methods often overlook these variant branches. Huang et al. [12] treated abnormal bronchus as a separate category, while the absence of anatomical standards made it challenging to categorize them equally with other classes as illustrated in Fig. 1(g).
Figure 1: Overview of bronchial anatomical labeling and its challenges. (a) and (b) illustrate the preoperative planning path for bronchoscopy guided by hierarchical anatomical labeling. (c) and (d) show successful and failed navigation trajectories, respectively. (e) compares node-wise correctness and subtree-wise consistency under correct labeling ($L_{C}$) and incorrect labeling ($L_{I}$). The results indicate that anatomical subtree-wise consistency is more sensitive to labeling errors. (f) and (g) demonstrate how individual variability and abnormal branches pose challenges to labeling accuracy and consistency
@@
To alleviate the aforementioned challenges, we focus on the topological consistency for bronchial anatomical labeling. Specifically, we propose the Soft Subtree Consistency (SSC) module to tackle the first challenge, SSC constructs a soft subtree via feature interaction and topological refinement. Segmental labels serve as the criterion for soft subtrees, resulting in more flexible and discriminative feature aggregation within and across subtrees. To address the second challenge, we propose an Abnormal Branch Saliency (ABS) module, which compares node features and prototypes to predict anomaly scores, distinguishing abnormal nodes. And a soft mask derived from anomaly scores is utilized to prevent feature aggregation between normal and abnormal nodes for a smoother regularization.
Building on a transformer-based framework, we introduce anatomical subtree-wise consistency as an evaluation standard to better reflect the integrity of bronchial structures. Compared to existing methods that focus on branch-wise labeling, our approach achieves notable improvements under this clinically relevant metric, particularly in complex or variant airway morphologies.
@@
The proposed method is evaluated on ATL-fibrosis dataset, a challenging dataset we constructed to facilitate comprehensive evaluation of airway labeling methods. Extensive experiments demonstrate that our method achieves superior performance compared to other state-of-the-art methods, achieving 91.4% accuracy in segmental, 83.7% accuracy in subsegmental with a 1.4% increase, 96.8% consistency with a 3.1% increase. The results demonstrate robustness across a diverse range of airway conditions, including severe abnormalities and anatomical variations, relevant to clinical scenarios.
@@ 2 Method
Figure 2: Overview of the proposed method: The airway labeling method comprises a Soft Subtree Consistency (SSC) module, which generates a soft subtree probability map, and an Abnormal Branch Saliency (ABS) module, which produces an abnormality score prediction. Both of them are used to inform attention interaction at the subsegmental level
Each airway branch is represented as a node in the graph, with attributes designed to reflect its location, morphology, and structure. An overview of the proposed method is illustrated in Fig. 2. We propose two modules to improve the consistency of prediction: (1) SSC module (detailed in Section 2.2) softly encodes the relationship between subtrees. (2) ABS module (detailed in Section 2.3) distinguishes between normal nodes and abnormal nodes.
@@ 2.1 Overall Framework
We adopt a U-shaped framework that progressively improves classification accuracy by incorporating hierarchical relationships. This framework comprises two stages: an independent stage ($1st$ stage) and a guided refinement stage ($2nd$ stage). The latter employs a fine-to-coarse guidance mechanism, using finer features from the independent stage to refine coarser predictions, thereby enhancing overall labeling performance.
@@
$\tilde{G}_{m}^{i}=\mathcal{F}(q=k=v=X_{m}^{i}),$ (1)
$G_{m}^{i}=\left\{\begin{array}[]{lll}\mathcal{F}(q=\tilde{G}_{m}^{i},k=v=G_{m+1}^{i-1})&&\text{for }m\in\{lob,seg\}\text{ and }i=2\\ \tilde{G}_{m}^{i}&&\text{otherwise}\end{array}\right.$ (4)
$X_{m+1}^{i}=G_{m}^{i}\quad\text{ for }m\in\{lob,seg\}$ (5)
$P_{m}^{i}=\mathcal{C}_{m}^{i}(G_{m}).$ (6)
Here, the hierarchical level $m\in\mathcal{H}$, where the ordered set =$\mathcal{H}=\{lob,seg,sub\}$ corresponding to lobar, segmental, and subsegmental levels respectively. $\mathcal{F}$ denotes the function of Transformer architecture. $X_{m}^{i}\in\mathbb{R}^{N\times d}$ and $G_{m}^{i}\in\mathbb{R}^{N\times d}$ represent the input and output features of the m-th transformer block in the i-th stage, where $N$ denotes the number of nodes and $d$ is the feature dimension. $\mathcal{C}_{m}^{i}$ denotes the m-th classifier consisting of a simple linear layer, while $P_{m}^{i}\in\mathbb{R}^{N\times c_{m}}$ is the prediction result where $c_{m}$ denotes the number of categories. For simplicity, we take the guided refinement stage for an example, omitting the index $i$ in subsequent notation.
@@ 2.2 Soft Subtree Consistency
Due to significant inter-individual variability, relying solely on node representations for classification can lead to substantial feature overlap. To address this challenge, we propose the SSC module, which dynamically encodes subtree representations, unlike hard-subtree techniques which rely on descendant counts for subtree construction.
We utilize segmental labels to define subtrees that convey clinically relevant anatomical meaning. The soft subtree map $M_{t}\in\mathbb{R}^{N\times N}$ is introduced, where $M_{t}(i,j)$ represents the probability that nodes $v_{i}$ and $v_{j}$ belong to the same subtree. The ground truth labels for $M_{t}$ are defined as a binary map that represents the probability of two nodes sharing the same segmental label. Specially, $M(i,j)$ is 1 if node $v_{j}$ is a descendant of node $v_{i}$, and 0 otherwise.
Soft subtree map $\bar{M}_{t}$ is inferred through segmental feature $G_{seg}=(g_{1},g_{2},\dots,g_{N})^{T}\in\mathbb{R}^{N\times d}$, where $g_{i}$ represents the feature of node $v_{i}$. The inference involves each node feature $g_{i}$ interacting globally with all other nodes to capture intra-subtree relationships, followed by dimensionality reduction using learnable $1\times 1$ convolutional layers(as shown in Fig. 2 (a)). Then referencing the hard subtree which consists of a node and its descendants, the probability of two nodes belonging to the same subtree is increased if one node is a descendant of the other. To achieve this, a soft topological refinement is applied, where descendant relationships $M_{D}$ dynamically adjust the probabilities in $M_{t}$:
$\hat{M}_{t}=M_{D}\odot(\bar{M}_{t}+K\odot(1-\bar{M}_{t}))+(1-M_{D})\odot\bar{M}_{t},$ (8)
where $K\in[0,1]^{N\times N}$ is obtained from $G_{seg}$ to modulate the refinement. $M_{D}\in\mathbb{R}^{N\times N}$ is a binary mask that indicates the descendant relationship between nodes. Specifically, $M_{D}(i,j)$ is 1 if node $v_{j}$ is a descendant of node $v_{i}$, and 0 otherwise.
@@
According to the hierarchical nomenclature, the assignment of subsegmental labels is determined by their relative relationships to branches within a segment. Therefore, the soft subtree map is utilized to constrain feature aggregation within and across subtrees at the subsegmental level as follows,
@@
Feature overlap occurs when node representations from different anatomical pattern become indistinguishable, particularly in cases with substantial morphological variability. The SSC module proves particularly effective in handling anatomical variations. For example, Fig. 3 (b-d) illustrates a scenario where the positions and orientations of LB4 branches in this variant differ significantly from the typical pattern (Fig. 3 (a)), resulting in substantial feature overlap between LB3 and LB4. The SSC module alleviates this issue by constraining feature aggregation within subtrees softly (as shown in Fig. 3 (e-f)), ensuring consistency even in cases of significant anatomical variability. Moreover, by integrating topological refinement with attention mechanisms, it not only improves subtree clustering but also enhances overall anatomical coherence.
@@ 2.3 Abnormal Branch Saliency
The position, orientation, and shape of abnormal branches are uncertain, with significant individual differences. The primary criterion for identifying an abnormal branch is its inability to be classified into any category within the nomenclature. Thus, we propose an ABS module that compares node features with class representations to detect anomalies. First, the prototypes are obtained through the segmental features $G_{seg}\in\mathbb{R}^{N\times d}$ and prediction result $P_{seg}\in\mathbb{R}^{N\times c_{seg}}$ :
$S_{seg}$ $=softmax(P_{seg}),$ (10)
$H$ $=((S_{seg})^{\alpha})^{T}G_{seg},$ (11)
$S_{seg}\in[0,1]^{N\times c_{seg}}$ denotes the probability matrix of the segmental level. $H\in\mathbb{R}^{c_{seg}\times d}$ are representations for segmental categories, while $\alpha$ is a learnable cluster parameter. Then they are refined through a transformer to incorporate global contextual information and ensure better class representations for each segmental category:
@@
After obtaining the refined class representations, node features interact with these prototypes $H$ to compute an anomaly score. As shown in Fig. 2 (b), each pair of feature and prototypes is concatenated along the depth dimension, forming the $\phi\in\mathbb{R}^{N\times c_{seg}\times 2d}$. $\phi$ is processed through $1\times 1$ convolutional layers to reduce the depth dimension, followed by the Multilayer Perceptron (MLP) to compute the anomaly score $\hat{Y}_{a}\in[0,1]^{N\times 1}$, indicating the probability of each node being an abnormal branch.
Due to the distinctive characteristics of abnormal branches, it is too absolute and arbitrary to use this anomaly score to directly classify nodes as a category equally with other classes. Instead, we generate a soft mask $M_{a}\in\mathbb{R}^{N\times N}$ for smoother regularization, which prevents feature aggregation between normal and abnormal nodes:
@@
Mask $\hat{M}_{a}$ ensures that abnormal nodes are guided away from normal nodes in the feature space, reducing their disruptive influence on normal branch predictions.
To further enhance subsegmental feature interactions, the combination between two soft masks $\hat{M}_{t}\odot\hat{M}_{a}$ , is applied to the attention interaction of subsegmental Transformer Block,
@@
The ABS module not only identifies abnormal branches but also reduces their disruptive influence on normal branch consistency, indirectly enhancing the overall accuracy and robustness of the model. Additionally, the anomaly score provides valuable insights for clinicians to identify structural abnormalities and understand airway variations, improving the practical utility of the method.
For classification tasks, cross-entropy loss with label smoothing is used, excluding abnormal nodes. Binary cross-entropy loss is employed for soft subtree map $\hat{M}_{t}$ prediction and abnormal score prediction $\hat{Y}_{a}$. The total loss function is formulated as:
$\begin{split}\mathcal{L}=\sum_{i=1}^{2}\gamma_{i}\Bigl{(}&\sum_{m\in\mathcal{H}}\alpha_{m}\mathcal{L}_{CE}(\hat{P}_{m}^{i},Y_{m})\\ &+\beta_{1}\mathcal{L}_{BCE}(\hat{M}_{t}^{i},M_{t})+\beta_{2}\mathcal{L}_{BCE}(\hat{Y}_{a}^{i},Y_{a})\Bigr{)}\end{split},$ (15)